# SPDFusion: An Infrared and Visible Image Fusion Network Based on a Non-Euclidean Representation of Riemannian Manifolds

### v2

@@ 4 Experiments and analyses
In this section, we introduce the fusion results and the analysis of our method. Initially, the SPDFusion experimental settings are presented, followed by ablation studies that reveal the effectiveness of the proposed fusion method and of the SPD attention module. The implementation details are described for repeatability. SPDFusion is benchmarked against leading methods in image fusion tasks, supported by qualitative and quantitative experimental analyses. Finally, the ablation experiments relating to our method are presented and discussed, and these ablation experiments were all conducted on the TNO dataset.
@@ 4.1.1 Datasets
In our work, we selected 1083 pairs of corresponding infrared and visible light images from the MSRS dataset [35] as training data, and some instances from the dataset are shown in Fig. 5. The size of the training images is standardized to 256×256 pixels. During the testing phase, we use 40 pairs of images from TNO [36] and 361 pairs of images from MSRS as the test sets, respectively. The dimensions of the test images are typically not fixed.
@@ 4.1.2 Parameter Setting
4.1.3 The Methods Compared and the Quality Metrics Used.
The method presented in this article was compared and evaluated with nine different state-of-the-art image fusion network approaches, including some classic and latest methods. These are: GANMcC [37], DenseFuse [38], FusionGAN [39], SemLA [40], FLFuse [41], RFN-Nest [14], DeFusion [42], DDFM [43] and LRRNet [9].Regarding the quality metrics, seven indices were chosen for performance evaluation, including: Entropy (En), Mutual Information (MI), Standard Deviation (SD), Spatial Frequency (SF), Visual Information Fidelity (VIF), Average Gradient (AG), $Q^{AB/F}$.These indices measure the overall quality of images from different perspectives.
@@ 4.2 Comparative Results Analysis
In this section, we conduct both qualitative and quantitative experiments with the proposed SPDFusion using two classical infrared-visible light datasets, TNO and MSRS, to verify the performance of our method.
TABLE I: Quantitative Experiments on the TNO Dataset.
Methods Year EN MI SD SF VIF AG Qabf
DenseFuse 2019 6.380 2.227 25.259 6.570 0.574 2.569 0.349
FusionGAN 2019 6.551 2.330 30.416 6.306 0.413 2.422 0.230
GANMcC 2021 6.735 2.276 33.361 6.189 0.520 2.553 0.277
RFN-Nest 2021 6.971 2.126 37.090 5.931 0.550 2.691 0.333
DeFusion 2022 6.607 2.627 30.868 6.461 0.559 2.624 0.365
FLFuse 2023 6.388 2.177 25.600 6.767 0.603 2.622 0.395
SemLA 2023 6.879 2.051 37.890 8.528 0.470 2.892 0.265
LRRNet 2023 6.991 2.518 40.984 9.608 0.548 3.790 0.352
DDFM 2023 6.854 2.231 34.418 8.629 0.631 3.405 0.434
SPDFusion Ours 7.055 2.540 41.936 9.655 0.645 3.973 0.405
TABLE II: Quantitave Experiments on the MSRS Dataset.
Methods Years EN MI SD SF VIF AG Qabf
DenseFuse 2019 5.931 2.666 23.550 6.020 0.692 2.053 0.368
FusionGAN 2019 5.431 1.893 17.076 4.354 0.442 1.452 0.140
GANMcC 2021 6.120 2.565 26.356 5.664 0.635 2.006 0.302
RFN-Nest 2021 6.196 2.460 29.078 6.163 0.656 2.115 0.390
DeFusion 2022 6.383 2.990 35.429 8.146 0.730 2.654 0.507
FLFuse 2023 5.754 2.134 19.469 4.841 0.651 1.815 0.289
SemLA 2023 6.423 2.461 33.122 6.351 0.617 2.257 0.290
LRRNet 2023 6.192 2.928 31.758 8.473 0.541 2.651 0.454
DDFM 2023 6.175 2.735 28.925 7.388 0.743 2.522 0.474
SPDFusion Ours 6.362 2.786 35.746 8.928 0.740 2.878 0.524
@@ 4.2.1 The Fusion Results on TNO
First, we conduct experiments on the TNO dataset, by comparing all the methods on an image called ”$man$”, which is shown in Fig. 6.
@@
It is worth mentioning that for the parts of the image involving the tree tops and the sky, as indicated by the red highlighted areas, previous methods such as DenseFuse, FusionGAN, GANMcC, and FLFuse were unable to maintain the original color features and texture details of the images. In comparison, although RFN-Nest and SemLA provided some improvements in brightness and details in the visible light, their capacity to preserve essential details in edge areas, like pedestrian outlines, remained a problem. Meanwhile, the latest three methods, DeFusion, LRRNet, and DDFM, perform well in these aspects, but it is still difficult to achieve a balance between capturing high correlation scene information and obtaining low correlation semantic information in different modalities. Our SPDFusion enhanced the clarity and brightness of the image in areas where the background is more indistinct. On the other hand, in the parts concerning the targeted individuals, as shown by the yellow highlighted areas, our approach was able to offer clearer delineation of the edge regions, resulting in higher contrast, while also conforming to human visual perception.
@@
Additionally, we utilized the previous seven indices to conduct an objective evaluation of the methods mentioned above. As shown in the Table I, the first and the second optimal results are marked with bold and red font respectively. Among the various state-of-the-art fusion methods, SPDFusion achieved the highest level in five classic image fusion metrics (EN, SD, SF, VIF, AG) and reached the second-highest level in the remaining two metrics (MI, $Q^{AB/F}$). These results indicate that our method not only enhances the overall sharpness of the image but also presents a more balanced and realistic visual effect, significantly surpassing existing technologies in terms of visual quality.
4.2.2 Fusion Results on MSRS
To verify the generalization performance of our SPDFusion, we selected more IR-VI pairs from the MSRS dataset for analysis and comparison, as shown in Fig. 7. Compared to methods like FusionGAN, GANMcC, RFN-Nest, SemLA and DeFusion, our SPDFusion method highlights infrared targets and performs visual enhancement more distinctively, as illustrated by the yellow boxed sections in the images. This to some extent reflects that the Riemannian attention mechanism can capture deep statistical information across different modalities effectively. At the same time, fusion images generated by DenseFuse, FLFuse, LRRNet and DDFM appear relatively blurry, whereas our method retains more detailed parts of the objects within the scene (as indicated by the red boxed sections in the images). This means that our method preserves not only the prominent features from the infrared image but also the texture details from the visible light image, which is crucial for the task of image fusion.
In the Table II, compared to other fusion methods, our proposed SPDFusion achieved the best values in four metrics (SD, SF, AG, $Q^{AB/F}$) and the second-best in one metric (VIF). This indicates that the images fused via SPDFusion contain more edge details and have a higher degree of information fidelity, also reflecting our work’s success in achieving cross-modal pixel and semantic-level fusion.
4.3 Ablation Study
In this section, we objectively analyze the effectiveness of several modules in our work and conduct a comparative analysis addressing the following aspects: the fusion strategies, different layer structures of SPD-block, the rationality of manifold attention and selection of different $\epsilon$ values.
4.3.1 Different fusion strategies
In this section, we discussed in detail how different fusion strategies affect the fusion results and compared them from subjective and objective perspectives. ”cross-modal strategy” refers to a fusion strategy that computes the covariance matrices of two modalities, while ”single-modal strategy” refers to computing the covariance matrices of each modality individually. Table III presents the average values of five metrics, with the best values shown in bold. From a quantitative perspective, the cross-modal fusion strategy exhibits better performance in all five metrics: EN, SD, SF, VIF, and AG. This demonstrates that the design of our fusion strategy is beneficial in making the fused images visually closer to real scenes, effectively preserving the rich texture details of the source images. Therefore, in our framework, we adopt a cross-modal fusion strategy to integrate information from multiple source images and generate high-definition fused images. Some examples of the fusion results obtained using different fusion strategies are shown in Fig. 8.
IR/VI
Cov(ir+vi)
Cov(ir)
Cov(vi)
Figure 8: The qualitative results obtained using unimodal and multimodal fusion strategies.
4.3.2 Different Layer Structures of SPD-Block
Deep architectures may not always exhibit optimal performance in low-level visual tasks such as image fusion. Therefore, we attempted to train using different manifold network structures. As shown in Table III and Fig. 9, when the number of layers of BiMap and ReEig is only set to 1, the features of the original data are faithfully preserved through single-layer transformation, reducing the distortion or information loss that may occur in the features after multi-layer mapping. Meanwhile, as the number of layers in the SPD network increases, the gradient update may bias towards the deep semantic features within the modality due to their larger variance, while fusion tasks often require more attention to pixel-level information. Therefore, in our work, we adopt a layer of BiMap and ReEig to achieve the best performance.
@@ 4.3.2 Different Layer Structures of SPD-Block
DenseBlock
SEBlock
ViT
CBAM
SPDAM
Figure 10: Infrared and visible image fusion experiment on the TNO dataset.
@@
Figure 11: Visualization of the impact of different $\epsilon$ values on the fusion results.
@@
TABLE III: Objective results of ablation study. Here, ”b2-r2” and ”b3-r3” represent the number of blocks in BiMap and ReEig, respectively. ”single-modal strategy(ir)” and ”single-modal strategy(vi)” indicate the modal information contained within the SPD matrix that needs to be learned. ”$F_{Dense}$”, ”$F_{SE}$”, ”$F_{CBAM}$”, ”$F_{ViT}$” denote that our SPDAM is replaced by Dense Block, SE Block, CBAM, and ViT structures, respectively. ”$\epsilon$” is a perturbation applied to the images.
EN SD SF VIF AG
Ours(b1-r1) 7.055 41.936 9.655 0.645 3.973
b2-r2 6.995 38.224 8.762 0.552 3.761
b3-r3 6.979 37.575 8.758 0.550 3.788
single-modal strategy(ir) 6.951 36.343 8.583 0.529 3.727
single-modal strategy(vi) 7.024 38.733 8.564 0.527 3.675
$F_{Dense}$ 6.914 35.626 9.072 0.592 3.987
$F_{SE}$ 7.003 38.893 7.954 0.520 3.520
$F_{CBAM}$ 6.927 36.296 6.524 0.506 2.827
$F_{ViT}$ 6.906 37.356 9.662 0.633 3.799
$\epsilon$=0.01 6.969 37.749 9.080 0.589 3.871
$\epsilon$=0.0001 7.010 36.917 8.319 0.406 3.637
$\epsilon$=0.005 6.975 37.896 9.075 0.583 3.883
$\epsilon$=0.0005 6.976 37.028 8.707 0.511 3.784
@@ 4.3.3 Manifold Attention
In this section, we assess the effectiveness of SPDAM by using different network architectures, as illustrated in Table III. We replaced our attention module with four of the most popular network structures currently available, including the relatively classic network architecture (DenseNet) and attention-based network architectures (SENet, CBAM, Transformer). For a more intuitive comparison, we visualize both the source images and the fused images from different methods. As shown in Fig. 10, compared to previous traditional methods, our SPDFusion enhances the texture features of shrubbery and trees and also highlights the detail features of the pedestrian area in the infrared modality. In addition, in our example, highly relevant parts like the sky have been significantly enhanced.
@@
At the same time, we also conducted a quantitative comparative experiment to further demonstrate the effectiveness of our SPDAM in the fusion task. As shown in Table III, our module surpasses other methods in terms of EN, SD, and VIF indicators, indicating that under the iteration of the Riemannian network, the network has deepened its focus on the associated information of the salient parts within modalities. Moreover, it emphasizes the useful complementary information between different modalities, promoting fusion with high contrast and detailed preservation.
4.3.4 Setting the Size of $\epsilon$
Equation 3 enhances the discriminative of SPD matrices through the correction of small eigenvalues. Since the shortest path between points on a Riemannian manifold is represented by geodesic distance, the epsilon setting in the ReEig layer is effectively adjusting the curvature on the manifold, thereby affecting the representation of local contrast and textural information in images. If $\epsilon$ is too large, it may cause data points to shift excessively on the manifold, resulting in the loss of the original geometric structure. Conversely, if $\epsilon$ is too small, it may lead to certain edge or detail features not being emphasized, making it difficult for the fused image to capture important information from different modalities. Therefore, we conduct an ablation study to determine the effect of different values of $\epsilon$. Their values are set as follows: $\epsilon\in\left\{1\mathrm{e}^{-2},5\mathrm{e}^{-3},1\mathrm{e}^{-3},5\mathrm{e}^{-4},1\mathrm{e}^{-4}\right\}$.
As shown in Fig. 11, we selected an instance from the MSRS dataset for intuitive demonstration. When $\epsilon$ is set to $1\mathrm{e}^{-2}$ and $1\mathrm{e}^{-4}$, the image exhibits obvious block artifacts due to uneven SPD weight distribution, which affects the visual quality of the image. When $\epsilon$ is set to $5\mathrm{e}^{-3}$ and $5\mathrm{e}^{-4}$, the left part of the image has many details submerged in darkness, failing to provide effective scene brightness information. When $\epsilon$ is set to $1\mathrm{e}^{-3}$, the fused image achieves the optimal effect.
Therefore, in the subsequent experiments, we set our $\epsilon$ to $1\mathrm{e}^{-2}$.
(a) DenseFuse
(b) FusionGAN
(c) GANMcC
(d) RFN-Nest
(e) DeFusion
(f) FLFuse
(g) SemLA
(h) LRRNet
(i) DDFM
(j) SPDFusion
Figure 12: Object detection results compared with other state-of-the-art fusion methods.

### v3

@@ 4 Experiments and Analyses
This section presents fusion results and systematically evaluates SMLNet. We first detail the experimental setup, then conduct ablation studies to validate the effectiveness of both the SPD attention module and fusion strategy. Comprehensive benchmarks on image fusion tasks and downstream applications, supported by qualitative and quantitative analyses, demonstrate the superiority of our method. Efficiency comparisons further highlight the advantages of our lightweight design. To ensure a thorough assessment, we also perform analysis of the method’s performance under extreme scenarios.
@@ 4.1.1 Datasets
In our work, we selected 1083 pairs of corresponding infrared and visible images from the MSRS dataset  (Tang et al., 2022b) as training data, ensuring a diverse representation of multi-modal scenarios. The size of the training images is standardized to 256×256 pixels. During the testing phase, we use 40 pairs of images from TNO  (Toet and Hogervorst, 2012), 361 pairs of images from MSRS and 40 pairs of images from RoadScene (Xu et al., 2020) as the test sets, respectively. The dimensions of the test images are typically not fixed, and some instances from the dataset are shown in Fig. 5.
@@ 4.1.2 Parameter Setting
4.1.3 The Methods Compared and the Quality Metrics Used
The method presented in this article was compared and evaluated with thirteen different state-of-the-art image fusion network approaches, including some classic and latest methods. These are: DenseFuse (Li and Wu, 2018), RFN-Nest (Li et al., 2021a), DeFusion (Liang et al., 2022), SemLA (Xie et al., 2023), DDFM (Zhao et al., 2023b) , LRRNet (Li et al., 2023), IRFS (Wang et al., 2023b), CrossFuse (Li and Wu, 2024), VDMUFusion (Shi et al., 2024), DCINN (Wang et al., 2024), DDBFusion (Zhang et al., 2025), FusionBooster (Cheng et al., 2025b) and GIFNet (Cheng et al., 2025a). Regarding the quality metrics, seven indices were chosen for performance evaluation, including: Entropy (En), Mutual Information (MI), Standard Deviation (SD), Spatial Frequency (SF), Visual Information Fidelity (VIF), Average Gradient (AG), $Q^{AB/F}$.These indices measure the overall quality of images from different perspectives, and the descriptions of the metrics can be found in (Ma et al., 2019).
@@ 4.2 Comparative Results Analysis
In this section, we carry out qualitative and quantitative experiments with the proposed SMLNet on the aforementioned three classical infrared-visible datasets to verify the method performance.
Figure 9: The qualitative results obtained using single-modal and multi-modal fusion strategies. When the second-order statistical information of the infrared modality is disregarded, the fusion results appear darker with lower contrast. Conversely, excluding the visible modality from covariance computation leads to blurred outcomes.
@@ 4.2.1 The Fusion Results on TNO
First, we conduct experiments on the TNO dataset, which are shown in Fig. 6.
@@
It is worth mentioning that for the parts of the image involving the tree tops and the sky, as indicated by the yellow highlighted areas, methods such as (c), (d), (f) and (k) are unable to maintain the original color features and texture details of the images. In comparison, although (e), (i), (m), (n) and (o) have improvements in terms of visible light brightness and detail, their ability to preserve fundamental objection contours remains an issue, and they also introduce relatively unfriendly overexposure. Meanwhile, (g), (h), (j) and (l), perform well in these aspects, but it is still difficult to achieve a balance between capturing high correlation scene information and obtaining low correlation semantic information in different modalities. Our SMLNet enhanced the clarity and brightness of the image in areas where the background is more indistinct. On the other hand, in the parts concerning the targeted individuals, as shown by the red highlighted areas, our approach was able to offer clearer delineation of the edge regions, resulting in higher contrast, while also conforming to human visual perception.
@@
Additionally, we utilized the previous seven indices to conduct an objective evaluation of the methods mentioned above, as shown in the Table 1. Among the various state-of-the-art fusion methods, SMLNet achieved the highest level in five classic image fusion metrics (EN, MI, SD, VIF, $Q^{AB/F}$) and reached the second-highest level in the remaining two metric (SF and AG). These results indicate that our method not only enhances the overall sharpness of the image but also presents a more balanced and realistic visual effect, significantly surpassing existing technologies in terms of visual quality.
@@
4.2.2 Fusion Results on MSRS
To verify the generalization performance of our SMLNet, we selected more IR-VI pairs from the MSRS dataset for analysis and comparison, as shown in Fig. 7. Compared to methods like (d), (e), (f), (i), (l), (n) and (o), our SMLNet method highlights infrared targets and performs visual enhancement more distinctively, as illustrated by the red boxed sections in the images. This, to some extent, reflects that the Riemannian attention mechanism can effectively capture deep semantic information in the infrared modality. At the same time, fused images generated by (c), (g), (h), (j), (k) and (m) appear relatively blurry, whereas our method retains more detailed parts of the objects within the scene (as indicated by the yellow boxed sections in the images). This means that our method preserves not only the prominent features from the infrared image but also the texture details from the visible light image, which is crucial for the task of image fusion.
In the Table 2, compared to other fusion methods, our proposed SMLNet achieved the best values in four metrics (MI, SD, VIF and $Q^{AB/F}$) competitive results in two others (EN and SF). This indicates that the images fused via SMLNet contain more edge details and have a higher degree of information fidelity, also reflecting our work’s success in achieving cross-modal pixel and semantic-level fusion.
Figure 11: Conventional attention networks enforce modal alignment in Euclidean space while neglecting the manifold characteristics of cross-modal data, resulting in unnatural fused images with severe detail loss, and this often produces suboptimal fusion results. In contrast, our SPDAM adheres to the intrinsic topological structure between modalities through manifold space modeling, achieving a balance between detail preservation and semantic integrity.
4.2.3 Fusion Results on RoadScene
Fig. 8 presents a qualitative comparison on the RoadScene dataset. (c), (f), (g), (k) and (m) have lost some critical road marking information, which is quite crucial in the overall image. In the results obtained by (e), (i), (n) and (o), the pedestrians and ground appear too bright, which is not friendly to the human visual perception. (d), (h), (j) and (l) generally produced color distortion in the targets, leading to the loss of texture details. In contrast, the proposed method maximizes the retention of scene details while highlighting the intensity information of pedestrians in the infrared modality, achieving perfect fusion.
The average values of the seven metrics are shown in the Table 3, where SMLNet has reached the highest level on four metrics, even demonstrating remarkable advantages on three of them (MI, SD, VIF). The aforementioned observations indicate that the proposed manifold attention model can skillfully handle various complex scenes, producing clearer and more natural fusion results.
@@ 4.2.3 Fusion Results on RoadScene
w/o orth
Figure 12: Visualization of the impact of different $\epsilon$ values on the fusion results. Optimal SPD matrix regularization (e.g., $1\mathrm{e}^{-3}$) balances detail preservation and semantic integrity, while improper values cause artifacts or loss of critical features. In addition, the BiMap layer’s orthogonality constraint is essential to maintain Riemannian geometry, as random initialization distorts manifold structure and degrades fusion quality.
@@
Table 4: Objective results of ablation study on TNO dataset. Here, “b2-r2” and “b3-r3” represent the number of blocks in BiMap and ReEig, respectively. “single-modal strategy(ir)” and “single-modal strategy(vi)” indicate the modal information contained within the SPD matrix that needs to be learned. $F_{\text{Dense}}$”, $F_{\text{SE}}$”, $F_{\text{CBAM}}$”, $F_{\text{ViT}}$” denote that our SPDAM is replaced by Dense Block, SE Block, CBAM, and ViT structures, respectively. $\epsilon$” is a perturbation applied to the images. The highest result is highlighted in bolditalic.
EN SD SF VIF $Q^{AB/F}$
Ours(b1-r1) 7.068 43.209 11.098 0.832 0.516
b2-r2 7.000 39.857 9.993 0.638 0.449
b3-r3 6.991 39.412 9.784 0.633 0.443
single-modal strategy(ir) 7.018 41.481 9.049 0.616 0.426
single-modal strategy(vi) 6.989 41.003 10.363 0.641 0.449
$F_{\text{Dense}}$ 6.902 35.517 7.322 0.579 0.395
$F_{\text{SE}}$ 7.102 46.090 6.337 0.564 0.302
$F_{\text{CBAM}}$ 7.121 49.124 7.625 0.673 0.378
$F_{\text{ViT}}$ 6.950 39.112 9.069 0.643 0.428
$\epsilon$=0.01 6.988 39.708 10.164 0.662 0.464
$\epsilon$=0.0001 6.974 37.939 8.929 0.504 0.381
$\epsilon$=0.005 7.001 39.863 10.054 0.652 0.456
$\epsilon$=0.0005 7.003 40.429 9.600 0.564 0.401
w/o orth 6.947 39.523 9.789 0.772 0.499
4.3 Ablation Study
In this section, we objectively evaluate the effectiveness of key components in our framework through comparative analyses, including fusion strategies, SPD-block layer structures, the rationality of manifold attention, different manifold constraints, manifold distributions, and noise robustness tests.
Figure 13: Covariance row vectors visualized via t-SNE under PCA versus SPD manifold learning dimensionality reduction, color-coded by quadrant and supplemented with fusion results. Traditional methods exhibit modality-biased clustering, manifesting as distinct separations between Fused Q1/Q2 (modality-1-dominant) and Q3/Q4 (modality-2-dominant) distributions, as demarcated by red guidelines. In contrast, our manifold-based framework establishes a unified statistical representation for cross-modal images, achieving homogeneous fusion of multi-modal statistics. This uniform distribution eliminates modality dominance artifacts while preserving thermally salient structures and textural details, directly enhancing fusion quality.
4.3.1 Different Fusion Strategies
In this section, we discussed in detail how different fusion strategies affect the fusion results and compared them from subjective and objective perspectives. “cross-modal strategy” refers to a fusion strategy that computes the covariance matrices of two modalities, while “single-modal strategy” refers to computing the covariance matrices of each modality individually. Table 4 presents the average values of five metrics. From a quantitative perspective, the cross-modal fusion strategy exhibits better performance in all five metrics: EN, SD, SF, VIF, and $Q^{AB/F}$. This demonstrates that the design of our fusion strategy is beneficial in making the fused images visually closer to real scenes, effectively preserving the rich texture details of the source images. Some examples of the fusion results obtained using different fusion strategies are shown in Fig. 9. When only the infrared (IR) modality is used, the results lose texture details from visible images. Conversely, relying solely on the visible (VI) modality causes the loss of target brightness information, which is unacceptable for image fusion tasks. Therefore, in our framework, we adopt a cross-modal fusion strategy to integrate information from multiple source images and generate satisfactory fused images.
Table 5: The average values of different manifold distribution metrics on the MSRS dataset. The top two metrics are highlighted in bolditalic and bold.
CM-NNR $S$ IMDR
PCA 0.375 0.045 1.052
SPD(Ours) 0.442 0.016 1.012
4.3.2 Different Layer Structures of SPD-Block
Deep architectures may not always exhibit optimal performance in low-level visual tasks such as image fusion. Therefore, we attempted to train using different manifold network structures. As shown in Table 4 and Fig. 10, when the number of layers of BiMap and ReEig is only set to 1, the features of the original data are faithfully preserved through single-layer transformation, reducing the distortion or information loss that may occur in the features after multi-layer mapping. Meanwhile, as the number of layers in the SPD network increases, the gradient update may bias towards the deep semantic features within the modality due to their larger variance, while fusion tasks often require more attention to pixel-level information. Therefore, in our work, we adopt one layer of BiMap and ReEig to achieve the best performance.
Figure 14: Euclidean networks employ linear operations (e.g., convolution, matrix multiplication) and demonstrate heightened sensitivity to both Gaussian and salt-and-pepper noise due to their lack of geometric constraints, resulting in significant feature map degradation. In contrast, manifold networks effectively suppress high-amplitude noise interference through nonlinear geometric mappings while preserving the structural integrity of critical features.
@@ 4.3.3 Manifold Attention
To evaluate the effectiveness of SPDAM, we conduct experiments by replacing it with four representative network architectures, including both classic designs (e.g., DenseNet) and attention-based structures (e.g., SENet, CBAM, Transformer). For a more intuitive comparison, we visualize both the source images and the fused images from different methods. As shown in Fig. 11, compared to previous traditional methods, our SMLNet enhances the color and brightness characteristics and also highlights the detail features of the pedestrian area in the infrared modality. This shows that the manifold attention follows the intrinsic topological structure between modalities, making the input end semantically tend to be balanced.
@@
At the same time, we also conducted a quantitative comparative experiment to further demonstrate the effectiveness of our SPDAM in the fusion task. As shown in Table 4, our module surpasses other methods in terms of SF, VIF and $Q^{AB/F}$, indicating that under the iteration of the Riemannian network, the network has deepened its focus on the associated information of the salient parts within modalities. Moreover, it emphasizes the useful complementary information between different modalities, optimizing fused results for enhanced contrast while preserving structural detail.
Table 6: The average quantitative results of different network noise additions on the TNO dataset. The top two metrics are highlighted in bolditalic and bold.
Gaussian EN SD SF VIF $Q^{AB/F}$
$F_{\text{Dense}}$ 6.943 36.238 7.404 0.459 0.284
$F_{\text{SE}}$ 7.048 40.969 5.453 0.453 0.240
$F_{\text{CBAM}}$ 7.033 40.357 6.837 0.472 0.274
$F_{\text{ViT}}$ 6.992 39.171 7.512 0.459 0.295
SMLNet(Ours) 7.070 42.199 7.830 0.469 0.292
SP EN SD SF VIF $Q^{AB/F}$
$F_{\text{Dense}}$ 6.899 34.950 8.925 0.432 0.310
$F_{\text{SE}}$ 7.022 40.514 7.692 0.498 0.308
$F_{\text{CBAM}}$ 6.989 40.614 9.739 0.443 0.296
$F_{\text{ViT}}$ 7.001 39.029 16.658 0.331 0.343
SMLNet(Ours) 7.044 40.736 11.482 0.487 0.369
4.3.4 The Impact of Different Manifold Constraints
Equation 3 enhances the discriminative of SPD matrices through the correction of small eigenvalues. Since the shortest path between points on a Riemannian manifold is represented by geodesic distance, the epsilon setting in the ReEig layer is effectively adjusting the curvature on the manifold, thereby affecting the representation of local contrast and textural information in images. As shown in Fig. 12, we selected an instance from the MSRS dataset for intuitive demonstration. When $\epsilon$ is set to $1\mathrm{e}^{-2}$ and $1\mathrm{e}^{-4}$, the image exhibits obvious block artifacts due to uneven SPD weight distribution, which affects the visual quality of the image. When $\epsilon$ is set to $5\mathrm{e}^{-3}$ and $5\mathrm{e}^{-4}$, the image loses edge details and shows diminished object contours. When $\epsilon$ is set to $1\mathrm{e}^{-3}$, the fused image achieves the optimal effect.
Meanwhile, as shown in Table 4, when $\epsilon$ is too large, the excessive diagonal perturbation oversmooths the covariance structure, suppressing subtle details and high-frequency components. This reduces EN and SD, as the fused image loses nuanced variations. Conversely, an overly small $\epsilon$ inadequately regularizes the covariance matrix, weakening the emphasis on cross-modal edge and texture correlations. This results in diminished SF, VIF, and $Q^{AB/F}$, as critical inter-modal features are insufficiently integrated. Therefore, by selecting an appropriate perturbation value, we achieve an optimal balance between preserving high-frequency image details and effectively regularizing the covariance matrix.
Moreover, the BiMap layer, by design, enforces orthogonality or near-orthogonality conditions on its weight matrices to ensure that transformations respect the underlying Riemannian geometry of the data. This is critical for maintaining the discriminative properties of the features, because it preserves angles, distances, and relative magnitudes during linear transformations.
When remove this constraint and initialize the matrix randomly, the transformation no longer adheres to the Riemannian structure. The random matrix may introduce arbitrary scaling, skewing, or even collapsing of the feature space. For example, the white highlights might collapse to lower values because the transformation no longer preserves the relative ordering. As shown in Fig. 12, a “distortion” occurs on the manifold, disrupting the semantic and geometric relationships that the network was originally designed to encode.
The observed results align with our theoretical expectation that removing the orthogonality constraint would lead to severe image degradation.
Figure 15: Object detection results compared with other state-of-the-art fusion methods. Unlike traditional methods that simply concatenate multi-modal features, SMLNet’s manifold fusion mechanism adaptively learns shared feature representations of key targets (e.g., edges and textures). This approach notably enhances the detector’s sensitivity to small and blurred objects while effectively reducing false positives and missed detections.
Figure 16: Segmentation results compared with other state-of-the-art fusion methods. Existing methods commonly suffer from semantic information loss and misclassification issues. In contrast, SMLNet employs a correlation-based semantic weighting mechanism to enhance the contrast of discriminative regions while effectively suppressing irrelevant background interference.